1. Calculate the change in the internal energy of the following systems and then determine whether the processes are endothermic or exothermic: (a) 45 kJ of heat is transferred from the system to the surroundings. Answer: AE- the process is (b) 605 J of work is performed by the surroundings to the system. Answer: AE (c) When 145 J of work is done by the surroundings to the system, 212 J of heat is transferred from the system to the surroundings. Answer: LAE = the process is the process is (d) A 20.0 g sample of water is heated from 25°C to 35°C, thereby absorbing 33.6 kJ of heat from the surroundings Answer: AE (e) An internal combustion engine performs 44 kJ of work to the surroundings and transfers 91 kJ of heat to the surroundings. Answer: AE = the process is the process is
Answer #1

a) 45 kJ Heat is transferred from system to surroundings

q = -45 kJ

ΔE is Negative the process is exothermic.

(An exothermic process is one that transfers heat to the surroundings.)

b) 605J of work is performed by the surroundings to the system

q = 605 J

ΔE = positive and the process is endothermic

c)

q = - 212 J

w = 145 J

∆E = q + w

∆E = -212 J + 145 J = - 67 J

The process is exothermic

d)

q = m c ΔT

m =20g

c is the specific heat capacity = 4.184 J⋅K−1g

ΔT is the temperature change. 350 C – 250C = 100 C = 283 K

q = 20g x 4.184 J⋅K−1g x 283 K

= 23681.44J =23.68 kJ

Absorb 33.6 kJ heat from surrounding

ΔE = positive and the process is endothermic